OpenStudy (anonymous):

for the function defined by the graph OAB, find the half-range Fourier sine series.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[f(x)={h\over a}x\qquad\qquad\qquad\forall\;0\le x\le a\\ f(x)={h\over 1-a}(1-x)\qquad\forall\;a\le x\le 1\\ \]